estion 20 of 25 > Assign formal charges to cach atom in the three resonance forms of N,O. N-N=0: :N=N- N=n=; Answer Bank
Answer #1

Comlin Page Formal charge an atom in a Lewis sto. I H total no. of I Total no. of non valence eol-bonding jit in the free lonDate N=N=0 Formal charge on nitrogen atom no 1. F. C = 5- 4 - 1 (4) = 5-4-2 = 5-6 = -1. Formal charge on nitrogen atom no. 2.